OpenStudy (anonymous):

Given the system of constraints, name all vertices of the feasible region. Then find the maximum value of the given objective function.

OpenStudy (anonymous):



hero (hero):

By the way, what is the given objective function?

OpenStudy (anonymous):

C=6x-4y I had the desmo calc open but I couldn't figure out how to use it haha.

OpenStudy (anonymous):

What do the bottom 4 mean?

hero (hero):

Those are just line segments I graphed that correspond with specific lengths along the x or y axis.

OpenStudy (anonymous):

I see, so I just use the points to find the maximum. I have another question I am almost finished with, can you check it?

hero (hero):

Yes, you can use the points to find the maximum. I can check your other work.

OpenStudy (anonymous):

k, give me a minute to finish it.

OpenStudy (anonymous):

This is the question: Use the elimination method. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 I added: -2x-y+z=-3 + 2x+3y+3z=5 and got 2y+4z=2 I got stuck here. I think I am supposed to multiply one of the equations so I can eliminate another variable, but Im not sure how

OpenStudy (ranga):

Using equations 2 and 3 you eliminated x. Use equations 1 and 2 and eliminate x again.

hero (hero):

I usually eliminate z variable.

OpenStudy (anonymous):

Don't I have to multiply one or both the equations by a number so I get additive inverses?

hero (hero):

That way, I'm left with a two equation system in terms of x and y.

hero (hero):

From there, it is easy.

hero (hero):

But in this case, it is probably wise to eliminate x variable.

OpenStudy (anonymous):

So, to do what ranga said, I have to multiply the 1st and 2nd equation by a number to get inverses? and then I can add them to eliminate

OpenStudy (ranga):

You are allowed to subtract equations too. You can subtract equation 2 from equation 1 to get rid of x.

OpenStudy (ranga):

Or think of it as multiplying equation 2 by -1 and adding it to equation 1.

OpenStudy (anonymous):

alright, let me do it and I'll tell you what I got.

OpenStudy (ranga):

okay, btw, the first one you got 2y+4z=2 can be simplified further by dividing by 2.

OpenStudy (anonymous):

I haven't learned to do that yet. I think I am just supposed to do it this way. I got 2x+1-z=3

OpenStudy (anonymous):

and now I am going to add the equations together

OpenStudy (ranga):

it should be 2x + y - z = 3

OpenStudy (anonymous):

That's what I meant haha sorry I got: 3y+2z=3

OpenStudy (ranga):

Yes. You got an earlier equation with just y and z and now you have a second one. You can eliminate z from both.

OpenStudy (anonymous):

So, do I add the two new equations and solve for z?

OpenStudy (ranga):

Write down your two equations here first.

OpenStudy (anonymous):

2y+4z=2 3y+2z=3

OpenStudy (ranga):

Yes. Multiply the second by -2 and add it to the first.

OpenStudy (anonymous):

When I multiplied the 2nd eq by -2 I got: -6y-4z=-6

OpenStudy (ranga):

Yes. Add it to first to get rid of z.

OpenStudy (anonymous):

I got -4y=-4 and simplified it to y=1

OpenStudy (ranga):

Yes. Put y = 1 in any one of the two equations and find z.

OpenStudy (anonymous):

into one of the original equations, right?

OpenStudy (ranga):

No the two we were just using that had only y and z variables.

OpenStudy (ranga):

2y+4z=2 or 3y+2z=3

OpenStudy (anonymous):

I got 4z=0 and if I use the other I get 2z=0 idk what I should do

OpenStudy (ranga):

Either one means z is 0. Zero multiplied by 2 or 4 is still 0.

OpenStudy (anonymous):

oh, ok. and then I plug in z into an original eq?

OpenStudy (ranga):

So you have y = 1, z = 0. Pick any one of the original 3 equations and substitute the values for y and z and then solve for x.

OpenStudy (anonymous):

I got x=1

OpenStudy (anonymous):

so the solution is (1,1,0)

OpenStudy (ranga):

Correct. I always put the values in all three original equations to see if it satisfies all three equations. This is one way to double check that we did not make any calculation mistake anywhere.

OpenStudy (anonymous):

k, let me check

OpenStudy (anonymous):

they are all correct. thank you so much.

OpenStudy (ranga):

you are welcome